OpenStudy (calculusxy):

Which fraction equals to -9/16? A. -3/4 B. 9/-16 C. 9/16 D. 3/4

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

any ideas?

OpenStudy (calculusxy):

I thought about the first one. But I am not too sure.

OpenStudy (phi):

negative divided by positive is negative. the answer has to be a negative

OpenStudy (anonymous):

It's going to be negative, for one

OpenStudy (phi):

that leaves out C and D

OpenStudy (calculusxy):

So if it is negative, then the sign cannot be in the denominator. It has to be in the numerator.

OpenStudy (anonymous):

No, -(1/2) = (-1/2) = (1/-2)

OpenStudy (phi):

the sign can be either place (but not both) notice -1/2 is - 1/2 and 1/-2 is - 1/2 all three ways mean the same number

OpenStudy (calculusxy):

That was what my teacher had said.

OpenStudy (anonymous):

It doesn't matter so long as at least one number is negative

OpenStudy (calculusxy):

So is it B?

OpenStudy (phi):

ask your teach what is 9 divided by -16 ?

OpenStudy (anonymous):

If you want to show a negative fraction, then yes, it's common practice to put the negative number on the top or to the side, but you don't have to

OpenStudy (phi):

or test with a calculator type in -9/16 and compare that answer to 9 divided by -16 you will get the same number

OpenStudy (calculusxy):

Wait.... So this is basically seeing what quotients match right?

OpenStudy (phi):

9 can be factored into 3*3 16 into 2*2*2*2 they have no common prime factors, so 9/16 cannot be simplified. so yes , for this problem top and bottom have to be 9 and 16 and the sign can be put in 3 different places: \[ \frac{-9}{16}, \frac{9}{-16} , - \frac{9}{16} \]

OpenStudy (calculusxy):

OKay thanks :) Gotta go bye

OpenStudy (phi):

of course, they could have done this - 18/32 (which simplifies to -9/16) to trick you